# Causality for Complex Continuous-time Functional Longitudinal Studies with Dynamic Treatment Regimes

### v1

@@ 3 Defining the Causal Quantity for Functional Longitudinal Data
This assumption ensures that the counterfactual outcome process under a dynamic treatment regime $\mathbb{G}$ is well-defined, regardless of the specific partitioning of the time interval. It guarantees that the sequence of counterfactual outcome distributions converges to a unique limiting distribution as the mesh of the partition approaches zero. This property is crucial for making meaningful causal inferences in the functional longitudinal data setting.
The parameter of interest is defined as
$\operatorname{\mathbb{E}}\{\nu(\bar{Y}_{\mathbb{G}})\},$ (53)
where $\nu$ is a user-specified bounded continuous function. This estimand includes those considering the marginal mean of outcomes under a STR in Ying (2024). Note that the definition of $\bar{Y}_{\mathbb{G},\Delta_{K}[0,\infty]}$ depends on the realization of $A^{\prime}$. However, the distributions of multiple versions of $\bar{Y}_{\mathbb{G},\Delta_{K}[0,\infty]}$ are the same. As (53) is on the expectation level, the parameter of interest is uniquely and well defined.
  • Though the data are allowed to be functional and the underlying data generating mechanism can have uncountably infinite number of treatment-confounder feedbacks, a finite-dimensional distribution intervention in Section 2.1 can still be considered, for example, intervening dosage of certain drug hourly or daily;
  • Likewise, a marked point process measure in Section 2.2 represents intervening both dosage and frequency of usage for certain drugs;
4 Causal Identification for Functional Longitudinal Data
We define
The following assumptions, which extend and generalize those introduced in Ying (2022), are necessary for identifying the causal effect of interest in the context of dynamic treatment regimes:
Assumption 2 (Intervenable).
@@ 4 Causal Identification for Functional Longitudinal Data
Assumption 4 (Consistency).
For any discrete DTR $\mathbb{G}$, when $\bar{A}$ follows $\mathbb{G}$, $(\bar{Y},\bar{A},\bar{L},C)$ and $(\bar{Y}_{\mathbb{G}},\bar{A},\bar{L},C)$ are identically distributed.
@@
Assumption 5 (Continuous randomization).
@@
$\left\|\operatorname{\mathbb{E}}\{\nu(\bar{Y}_{\mathbb{G}})|\mathscr{F}_{t}\}-\operatorname{\mathbb{E}}\{\nu(\bar{Y}_{\mathbb{G}})|\bar{A}(t+\eta),\mathscr{F}_{t}\}\right\|_{1}<\varepsilon(t,\eta).$ (62)
@@
Furthermore, we need an assumption over the censoring mechanism to eliminate the censoring bias. We consider the conditionally independent censoring assumption in Van der Laan and Robins (2003); Rotnitzky and Robins (2005); Tsiatis (2006); Ying (2022). Define the full data censoring time hazard as
$\textup{d}\Lambda_{C}\{t|T,\overline{A}(T),\overline{L}(T)\}=\operatorname{\mathbb{P}}\{C\leq t+\textup{d}t|C>t,T,\overline{A}(T),\overline{L}(T)\}.$ (63)
Assumption 6 (Conditional independent censoring).
The censoring mechanism is said to be conditionally independent if
$\textup{d}\Lambda_{C}\{t|T,\overline{A}(T),\overline{L}(T)\}=\operatorname{\mathbb{P}}\{C\leq t+\textup{d}t|C>t,\mathscr{F}_{t}\}\mathbbm{1}(T>t).$ (64)
Definition 3 (G-computation process).
Under Assumptions 2 and 3, define
Theorem 1 (G-computation formula).
Under Assumptions 1 - 6, the marginal mean of transformed potential outcomes under a user-specified regime $\mathbb{G}$ (53) is identified via a g-computation formula as
Definition 4 (Inverse probability weighting process).
Under Assumptions 2 and 3, define
Theorem 2 (Inverse probability weighting formula).
Under Assumptions 1 - 6, the marginal mean of transformed potential outcomes under a user-specified regime $\mathbb{G}$ (53) is identified via an inverse probability weighting formula as
4.1 Recursive representation of g-computation process
@@
Under Assumptions 2, 3, and 13, for any $\mathscr{G}_{t}$-adapted process $Q(t)$ with $(H_{\mathbb{G}},Q)\in\mathcal{M}_{\textup{out}}$ and $\sup_{t}\|H_{\mathbb{G}}(t)Q(t)\|_{1}<\infty$, $\Xi_{\textup{g-comp},\mathbb{G}}(H_{\mathbb{G}},Q)$ is unbiased for zero, that is,
@@
4.2 Recursive representation of IPW process
@@
Under Assumptions 2, 3, and 13, for any $\mathscr{G}_{t}$-adapted process $H(t)$ with $(H,Q_{\mathbb{G}})\in\mathcal{M}_{\textup{trt}}$ and $\sup_{t}\|H(t)Q_{\mathbb{G}}(t)\|_{1}<\infty$, $\Xi_{\textup{IPW}}(H,Q_{\mathbb{G}})$ is unbiased for zero, that is,

### v2

@@ Appendix A Weak Assumptions
This assumption ensures that the counterfactual outcome process under a DTR $\mathbb{G}$ is well-defined, regardless of the specific partitioning of the time. It guarantees that the sequence of counterfactual outcome distributions converges to a unique limiting distribution as the mesh of the partition approaches zero. This property is crucial for making meaningful causal inferences in the FLD setting.
The following assumptions are weaker than the assumptions we gave but necessary for identifying the causal effect of interest in the context of dynamic treatment regimes:
Assumption A.2 (Intervenable).
@@
Assumption A.4 (Consistency).
$\bar{Y}=\bar{Y}_{\bar{A}}$ (32)
@@
Assumption A.5 (Sequential randomization).
@@
$\left\|\operatorname{\mathbb{E}}\{\nu(\bar{Y}_{\mathcal{A}})|\mathscr{F}_{t}\}-\operatorname{\mathbb{E}}\{\nu(\bar{Y}_{\mathcal{A}})|\bar{A}(t+\eta),\mathscr{F}_{t}\}\right\|_{1}<\varepsilon(t,\eta).$ (33)
@@
A.1 Other regularity conditions
Proposition 3.
Under Assumptions LABEL:assump:intercontinuity - 2,
whenever $|\Delta_{K}[0,\infty]|\to 0$. It is immediate that Assumption A.3 is equivalent to Assumption 3. Assumption A.2 holds and Assumption A.5 holds for any bounded function $\nu(\cdot)$.
Consequently, one might choose Assumptions 4 - 3 instead of Assumptions A.1 - 4 as starting point if “full conditional exchangeability” is more intuitive.
Assumption A.6 (Uniform $L^{2}$ convergence).
Proposition 4.
Under Assumptions A.6, the regularity assumptions considered in Theorem 5 holds, that is, the derivative and limit below can interchange,
Assumption A.7 (Observed TR Continuity).
[…]
See Ying (2022) for a complete discussion of this assumption.
Appendix B Recursive representation of two identification processes
B.1 Recursive representation of g-computation process
@@
Under Assumptions A.2, A.3, and .5, for any $\mathscr{G}_{t}$-adapted process $Q(t)$ with $(H_{\mathbb{G}},Q)\in\mathcal{M}_{\textup{out}}$ and $\sup_{t}\|H_{\mathbb{G}}(t)Q(t)\|_{1}<\infty$, $\Xi_{\textup{g-comp},\mathbb{G}}(H_{\mathbb{G}},Q)$ is unbiased for zero, that is,
@@
B.2 Recursive representation of IPW process
@@
Under Assumptions A.2, A.3, and .5, for any $\mathscr{G}_{t}$-adapted process $H(t)$ with $(H,Q_{\mathbb{G}})\in\mathcal{M}_{\textup{trt}}$ and $\sup_{t}\|H(t)Q_{\mathbb{G}}(t)\|_{1}<\infty$, $\Xi_{\textup{IPW}}(H,Q_{\mathbb{G}})$ is unbiased for zero, that is,